OpenStudy (anonymous):

f(x)=2x+2 solve f^-1(x) when x= 4 A. 1 B. 3 C. 4 D. 10

OpenStudy (anonymous):

f1 is 4

OpenStudy (anonymous):

f2 is 6

OpenStudy (anonymous):

f3 is 8

OpenStudy (anonymous):

f0 is 2

OpenStudy (anonymous):

f-1 is ?.?

OpenStudy (acxbox22):

find the inverse be swithching the x and y

OpenStudy (campbell_st):

start by swaping x and y so x = 2y + 2 now make y the subject lastly substitute x = 4 to find y

OpenStudy (acxbox22):

y=2x+2 x=2y+2 4=2y+2 2=2y y=1

OpenStudy (anonymous):

Thank you

OpenStudy (acxbox22):

\(\Huge\color{green}{\ddot\smile}\)

OpenStudy (acxbox22):

substitute -x+3 in the x for the function f(x)